### v1

@@ 6.1 Contributions
We introduce a novel sequence-to-sequence policy for generating an optimal trajectory in online planning scenarios, as well as a reinforcement learning algorithm that efficiently trains this policy by combining the evaluation from multiple sources. Specifically, the feasibility boundary is totally modeled from the data by training the reward estimator, which leads to generate a model that can fully utilize the vehicle’s capacity. In addition, the reward estimation module is trained with multi-fidelity Bayesian optimization, which allows us to efficiently incorporate real-world experiments into the optimization. The trained policy model generates faster trajectories with decent reliability compared to the baseline minimum snap method, even when the waypoints are randomly deviated and the baseline method fails.
6.2 Limitations
The main drawback of the proposed method is the absence of a performance guarantee. Although the trained model can generate faster and more feasible trajectories for the majority of the waypoint sequences, there are still some cases where the trained policy fails. This limitation could be addressed by training the policy with a more complex reward estimator model capable of capturing finer details of the feasibility boundary. Given that only the policy model is utilized during real-world deployment, the complexity of the reward estimator can be increased without requiring additional computational resources.
In addition, we have observed that the convergence time of reinforcement learning is influenced by the number of samples used for training the reward estimator. Increasing the number of labeled samples improves feasibility prediction accuracy, but it also requires more simulations and flight experiments. As inaccurate reward estimation increases the convergence time or often leads to suboptimal solutions, it is crucial to carefully tune the size of the labeled samples and balance the accuracy and computational requirements. By quantifying the discrepancy between different fidelity levels, we anticipate that convergence and optimality guarantees can be achieved.

### v2

@@ 6.1 Contributions
We introduce a novel sequence-to-sequence policy for generating an optimal trajectory in online planning scenarios, as well as a reinforcement learning algorithm that efficiently trains this policy by combining the evaluation from multiple sources. Our approach models the feasibility boundary entirely from data by training a reward estimator, enabling full utilization of the vehicle’s capacity. Furthermore, we employ multi-fidelity Bayesian optimization to train the reward estimation module, efficiently incorporating real-world experiments. This approach enables extensive policy optimization in regions where simulation fails to capture real-world phenomena. The trained policy model generates faster trajectories with decent reliability compared to the baseline minimum snap method, even when the waypoints are randomly deviated and the baseline method fails.
6.2 Limitations and Future Work
The main drawback of the proposed method is the absence of a performance guarantee. While the trained model generates faster and more feasible trajectories for most waypoint sequences, it occasionally fails. We expect this limitation could be addressed by leveraging the reward estimator beyond its primary role in training. The trained reward estimator can potentially predict output performance before physical testing, offering crucial guarantees for high-stake dynamic systems. This predictive capability could ensure robustness and safety in planners, and facilitate reachability analysis for safety-critical fields.
The long training time can be improved by redesigning the reward estimator’s architecture. We’ve observed that reinforcement learning convergence is influenced by the estimator’s accuracy. In this work, we use Gaussian process for feasibility prediction due to their effectiveness with limited samples. As data accumulates, however, a basic neural network becomes more suitable, offering better accuracy with larger datasets. A challenge arises when low-fidelity datasets reach this transition threshold earlier than high-fidelity ones, forcing continued use of Gaussian process and limiting low-fidelity sample inclusion. Developing a model that can transition between architectures and account for varying dataset sizes across fidelity levels could lead to more efficient large-scale multi-fidelity optimization, ultimately reducing training time.
Building upon the idea of adaptive architectures, increasing the reward model’s complexity and capacity could further extend the proposed method’s applicability. Since our approach treats the full system as a black box and optimizes tests based on specific goals, a more sophisticated model could incorporate additional system parameters. This enhanced capacity would allow the method to handle more complex elements, such as internal controller states, or integrate with local planners like MPC. Currently, this work only uses IMU and motion capture inputs, which could limit sources of perception uncertainty. A larger model capacity can ultimately extend this approach to full autonomy pipelines with various sensory inputs including camera and LiDAR, handling unique and high-dimensional uncertainties from the full system.